OpenStudy (anonymous):

Solve the following system by graphing. x - y = 4 x + y = 2 What is the solution of the system? (3, -1) (3, 1) (-1, 3)

pooja195 (pooja195):

Do you know where to start?

OpenStudy (anonymous):

nop

pooja195 (pooja195):

ok opposites cancel do you see opposite signs?

pooja195 (pooja195):

A positive and a negative?

undeadknight26 (undeadknight26):

Plug in your data! x - y = 4 3 = x -1 = y 3 - (-1) = 4?

OpenStudy (anonymous):

yes is 2 of them

pooja195 (pooja195):

ok opposites cancel so which varible cancels?